OpenStudy (inkyvoyd):

Differential equations - if you have a differential equation like y'' -2y'+2y=x, or something like that, what does it have to do with the solution of the corresponding polynomial?

OpenStudy (turingtest):

what is "it" ?

OpenStudy (turingtest):

...I don't think I understand your Q

OpenStudy (inkyvoyd):

i mean, it as in solution to differential equation. also, i'm just curious; just neefd a very very basic ubderstanding

OpenStudy (amistre64):

we assume there is a solution; and try to define a function that works in generality

OpenStudy (amistre64):

the details elude me; but e^r seems to be the best fit for these

OpenStudy (amistre64):

e^rx that as

OpenStudy (inkyvoyd):

HOw would go go about solving y'+2y''+1=e?

OpenStudy (amistre64):

is that a typo?

OpenStudy (turingtest):

easy second-order homogeneous eqn if not

OpenStudy (amistre64):

y = yh + yp

OpenStudy (amistre64):

assume e^rx is a solution to the homogenous part: y = e^rx y'= r e^rx y'' = r^2 r^rx y'' -2y' + 0y = 0 r^2 e^rx -2r e^rx +0 e^rx = 0 e^rx (r^2 -2r) = 0 ; when the poly goes zero

OpenStudy (amistre64):

r = 0 and r = 2

OpenStudy (inkyvoyd):

Ok, if I don't understand the assume e^(rx) is a solution ot he homogenous part, should I leave?

OpenStudy (amistre64):

yh = a e^0x + b e^2x

OpenStudy (amistre64):

lol, maybe :) the question assumes you know some baser things

OpenStudy (inkyvoyd):

Ok, if I want to learn, what are baser things?

OpenStudy (inkyvoyd):

also, what are the prerequisites of baser things?

OpenStudy (turingtest):

@inkyvoyd if you never heard of the characteristic polynomial or the stuff amistre is talking about I think you should start from the basics http://tutorial.math.lamar.edu/Classes/DE/DE.aspx

OpenStudy (amistre64):

yeah, pauls online is good stuff

OpenStudy (inkyvoyd):

Well, my 80's textbook on calc talks about it, and I was curious about the topic.

OpenStudy (amistre64):

i think the proofing comes from real analysis stuff

OpenStudy (turingtest):

prerequisites include: calculus I-III linear algebra (not strictly necessary, but I recommend it before learning DE's)

OpenStudy (inkyvoyd):

I have many holes in my calculus foundation, but I am really moreinterested in concepts that in a completely solid understanding, because I will eventually take calculus; right now I am just self studying

OpenStudy (inkyvoyd):

My txtbook is supposed to be calc 1-3?

OpenStudy (inkyvoyd):

And, linear algebra? Can I take that in parallel with calc?

OpenStudy (turingtest):

calculus is definitely before DE's you need a very solid calculus foundation before embarking on your study f DE's

OpenStudy (inkyvoyd):

(I know my comment about holes makes me sound like an idiot, and maybe I am, but I want to learn some of these harder things, because they interest me)

OpenStudy (turingtest):

study of*

OpenStudy (amistre64):

the best thing to do is to study up your foundational stuff in calc 1 thru 3 to have a foothold on whats going on in diffyQs i learned it up the same way

OpenStudy (inkyvoyd):

Oh, I know what I"m missing. I'm missing partial differentiation. Sorry.

OpenStudy (turingtest):

do you know integration by parts? or trigonometric substitution integrals?

OpenStudy (inkyvoyd):

I understand both.

OpenStudy (amistre64):

integration is the simplest diffyQ stuff that I know of

OpenStudy (inkyvoyd):

Won't say I would pass a test on them because I am very slow at calc (integration especially)

OpenStudy (inkyvoyd):

But I understand what they are

OpenStudy (inkyvoyd):

I'm not interested in being able to solve all of a certain type of differential equation, I just am curious about what's in the future for me.

OpenStudy (turingtest):

ok well then I suppose you could try your hand at DE's but that integration stuff is considered basic by the time you get to DE's, so you better practice that stuff 'till you can do it easily there are no shortcuts in math, unfortunately

OpenStudy (inkyvoyd):

So far my motivation to learn about math is pushed by my interest in complex numbers, but I'm sure it won't last until I actually get to learn about them (Complex analysis?)

OpenStudy (inkyvoyd):

So, I should be able to understand the link given to me?

OpenStudy (inkyvoyd):

(A basic understanding)

OpenStudy (turingtest):

not sure look at the whole page Paul's stuff goes from college algebra up through DE's so if you are confused, go back a few sections there is also a little section on complex numbers in one of the appendices

OpenStudy (inkyvoyd):

Thanks!

OpenStudy (inkyvoyd):

Sigh, I've been wondering about the applications of imaginary and complex numbers for a while now, and the closest I've gotten to even touching them is elecricity. My curiousity arose when I first learned about them and they were called "imaginary". Most people forget about them because they are completely "useless", but every textbook i've read say they have uses. Attempts to fin the applications led to hard math, and, the only thing I've learned about imaginary numbers that I can be proud of is euler's identity.

OpenStudy (inkyvoyd):

@TuringTest , scanned that page, and I think that's about as much as I need to know. I do realize that the people who developed solutions of differential equations were very good at calculus. :S Still though, I did understand what it was talking about for linear differntial equations, so I'll read on. Thanks again!

OpenStudy (turingtest):

very welcome :) be sure to thank Paul too, his notes are awesome !